OpenStudy (anonymous):

Assume that g(x,y) has an average value of g(avg)=2, Where R has a value of R(A)=3. Find the double integral of g(x,y)dA and (g(x,y)+1)dA

OpenStudy (anonymous):

check the question. the first sentence has something missing or mis-stated

OpenStudy (anonymous):

Preciously the question states Assume that g(x,y) has an average value of g(avg)=2 over the region R, Where R has an area of R(A)=3. Find the double integral of g(x,y)dA and (g(x,y)+1)dA

OpenStudy (anonymous):

"over the region R" -> missing in your original question also, does it say g(avg)?

OpenStudy (anonymous):

\[g_{avg}={1\over A}\iint g(x,y)dxdy\]

OpenStudy (anonymous):

Yeah it's saying the average value of g(x,y) is 2

OpenStudy (anonymous):

Right, and I see that we are given G(avg) and A but I don't know what to take the integrals from?

OpenStudy (anonymous):

region R has an area of "3" assume a square of side sqrt(3) \[0\le x,y\le\sqrt{3}\]

OpenStudy (anonymous):

So, both integrals are going from 0 to Square root of 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Cool, Thank you very much I appreciate it

OpenStudy (anonymous):

yep